If a red suit is drawn from an ordinary deck of cards, what is the probability that the card is a diamond?
Answer by oberobic(2304)

Given a red card is drown from the deck, then there are 26 total possibilities.
There are 13 ways the cared could be a diamond, which defines the favorable outcomes.
Probability = favorable outcomes / total possibilities
P(diamond) = 13/26 = 1/2